

********************************************************************
UARS WEEKLY STATUS REPORT - 20 NOV 1998
********************************************************************
********
SOLSTICE
********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occulation
data during this past week without apparent problems.
SUSIM was turned off during the Leonid storm. SUSIM was repowered
with no loss of daily data.
Reprocessing of all SUSIM experiment days to version 20 continues.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
HALOE STATUS
November 20, 1998
HALOE is powered off during this full graze period. HALOE is scheduled to
be powered on November 30, 1998.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
WINDII WEEKLY REPORT FOR THE WEEK ENDING NOV. 20, 1998
The WINDII Team has identified a long-standing problem with the
WINDII Doppler temperatures that has made them somewhat higher than
is considered to be correct. The correction involves a small change
in the software which is being implemented in the recently delivered
version.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF has received a total of 441 CD-ROM requests as of 20-November-1998.
There were four new orders received and shipped.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
21 days HAL (4.80) PROD (L2-L3) (62R-32C=30L)
179 days HAL (4.80) PROD (L3) (1874R-1349C=525L)
47 days SOL (4.62) PROD (L2-L3) (2313R-2064C=249L)
135 days SUS (7.00) PROD (L1) (2370R-404C=1966L)
6 days CLA (7.00) ENG (L1-L3) (18R-12C=6L)
CDHF pending software upgrades: IDL 5.1, OpenIngres, UCSS 10.1, MSS 4.1,
NRT 2.5, UOAS 5.4
***********
FOT
***********
REPORTING PERIOD
This report covers 07 November 1998 (Orbit 39103, GMT Day 311) through 14
November 1998 (Orbit 39208, GMT Day 317).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 311/0236 39105 On
WINDII 311/0611 39107 Off
WINDII 311/1259 39111 On
The beta angle ranged between 15.1 degrees to 45.5 degrees for this report
period. The beta angle is now increasing toward a maximum angle of 77.0
degrees on 22 November 1998 (DOY 326).
Flight Operations continues to work with the PEM instrument to fine tune
the operational command sequences associated with the OBC Power Monitor
day-only operational capability. Incremental progress this week shows
errors in both RTS and PDB activation of certain portions of the PEM Day-
only turn-on sequences. The problem is thought to be the result of command
buffer overflow with the increased pace of PDB and RTS commands.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.9 to 1.2 degrees C. Additional battery performance data
for each battery on 06 November 1998 and 13 November 1998 are provided in
Appendix D.
The clock error ranged from 10.1 to -12.1 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 01
% OUT-OF-LIMITS = 06
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 05
% PEM 06
% HRDI 01
% WINDII 03
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 03
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/286 - 2287.496917
Transponder B: 98/300 - 2287.498980
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/316 show the observatory data loss to be
21 hours, 11 minutes, 52.564 seconds (an increase of 13.984 seconds since
last report period). This is a 0.0339 percent data loss which equals a
99.9661 percent data capture for the mission.
The increase of 13.984 seconds occurred over seven (7) days:
7 2.656 seconds on Day 310 due to IP hits, and hits in data;
7 0.832 seconds on Day 311 due to hits in data;
7 0.832 seconds on Day 312 due to hits in data;
7 4.960 seconds on Day 313 due to hits in data and insufficient overlap;
7 2.944 seconds on Day 314 due to hits in data (IP);
7 1.248 seconds on Day 315 due to hits in data (IP); and
7 0.512 seconds on Day 316 due to hits in data.
Telemetry and Command Computer (TAC) RcrashesS have diminished this report
period.
The FOT modified the CMS resident Relative Time Sequence 32 (RTS 32) to
facilitate proper PMON-controlled day-only operation of the PEM instrument.
An anomaly was identified within the ephemeris dump/compare process
resulting from three miscompares (locations). The ground image miscompared
values were decoded (bit pattern) revealing that the three locations (words
1061 - 1063) contained NBTR stop, speed, and record commands. The CMS
generated spacecraft ephemeris contains 1060 commands. This load stops
short of the last three words in the System Table on purpose - last three
words are control functions. Therefore the miscompared locations in
question are outside of the loaded area but should have remained unchanged
since last loaded and verified good. To eliminate the miscompares and
provide a temporary solution, FOT overwrote the ground image with the
spacecraft values for these three (3) locations. The problem has been
reported.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
Investigation continues into possible impacts of the Leonid meteorite
shower in November 1998. A plan is being prepared and the necessary UARS
commanding identified.
Investigating the feasibility of a new PMON action to change battery V/T
curves based on loading to mitigate battery risks in sun pointing safehold.
YEAR 2000 OPERABILITY SUPPORT
UARS Test and Training Simulator (UTTS) - No change.
User Planning System (UPS) - No change.
Multi Satellite Operations Control Center (MSOCC) - No change.
Command Management System (CMS) - The FOT performed further CMS Y2K this
week. The testing is now approximately 95 percent complete. The remainder
of the testing will be completed when the FOT receives the last of the
necessary Y2K planning aids from FDF.
The FOT is still waiting for a Y2K 8-week S/C Orbital Events file from FDF
which contains a SC_YAWMAN event. Once the FOT receives this file, the
last remaining Y2K test function (#12) of the 1999-to-2000 test case (#1)
can be completed
Flight Dynamics Facility (FDF) - see CMS paragraph.
IP NETWORK TESTING
Space Network (SN) - Operational.
NCC 98 system - Complete (during SCS-21 demo verification).
Deep Space Network (DSN) and Ground Spaceflight Tracking and Data Network
(GSTDN) - test conducted with the Canberra (46) station on 10 November
1998. Command was successful plus 1.0 kbps and 32 kbps telemetry; 1 kbps
and 32 kbps OBC dump data; 512 kbps telemetry dump data and postpass replay
successfully flowed. The IP capability is considered satisfactory.
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
FOT continues to prepare the Latch valve command sequences and to plan the
testing with the UTTS simulator in preparation for the simulation with
hardware.
UARS RE-ENGINEERING - TELEMETRY PROCESSING
Flight operations is providing assistance on both sustaining engineering
and re-engineering activities to incorporate needed features into the
Trending system. Testing of the Trending System v.1.1 release (CM-0009)
continued. More DR's and ER's are being generated. Offline systems
engineering requested the on-line team keep track of the frequency of Multi
Functional Work Station (MFWS) lock-ups and reboots.
MEETINGS
UARS FOT personnel and Dave Lorenz met on 12 November 1998 to discuss the
preparation for the Leonid encounter on 17 November 1998. A second follow-
up meeting was held with the UARS FOT, Dave Lorenz, and Dave Ward (NASA ACS
Engineer) on 12 November 1998 to finalize the plans for the Leonid
encounter. Earth Sensor 2 will be turned ON and the EPSH (earth pointing
safe hold) will be selected for the high risk time range. The FHST 2
(fixed head star tracker) will remain ON in normal operation.
UARS FOT personnel met with T. Gitland, GSFC 530.4, R. Welborn, and D.
Gilliland of Stanford Telecom on 12 November 1998 to provide input for a
Space Network (SN) Study Team Review Project.
The next battery telecon is scheduled for 24 November 1998.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is scheduled for scheduled for 13
December 1998.
UARS will pass through the Leonid meteor shower on 17 November 1998 at
19:43 UTC (+/- 3 hours).
The United States Naval Observatory (USNO) issued a Time Service
announcement advising that the International Earth Rotation Service (IERS)
will retard UTC time by 1.0 second at the end of December 1998. This means
the addition of one leap second as the last second of 1998.
The next annular solar eclipse will occur on 16 February 1999 over the
Indian Ocean, Australia and the South Pacific Ocean.
PERSONNEL CHANGES
Mark Feeney joined on-line operations as an Flight Director on 09 November
1998. MarkUs previous experience includes ECS, Globalstar, EUVE and
Landsat.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 13 November 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
39105 311/0236 98-431 WINDII Power ON
39107 311/0611 98-432 WINDII Power OFF, S/C Low Pwr
39107 311/0611 98-433 OBC Clk Adj fr 0.00 to 20.6 msec/day
39109 311/0945 98-434 PEM RAM Dump
39111 311/1259 98-435 WINDII Power ON
39128 312/1450 98-436 OBC Clk Adj fr 20.6 to 0.00 msec/day
39145 313/1935 98-437 PEM PEM PDB update
39146 313/2108 98-438 PEM PEM RAM dump
39160 314/1926 98-439 HRDI Microprocessor load
39171 315/1243 98-440 OBC Clk Adj fr 0.00 to 20.6 msec/day
39175 315/1843 98-441 PEM PEM PDB testing
39188 316/1756 98-442 PEM PEM PDB testing
39204 317/1802 98-443 OBC Clk Adj fr 20.6 to 0.00 msec/day
39205 317/1938 98-444 PEM PEM PDB testing
39205 317/1938 98-445 Grd image Correct ephem image errors
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 13 November 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
NBRECBI Y-HI NBTR B Curren 39106-108,112
HRPSEQBIMON Y-HI PSE QB Current Mon 39107-8
WIQBCURRENT Y-LO WINDI QB Current 39110
PEMBIAS00E R-HI Sensor 0 Grid Bias 39188
PEMBIAS02E R-HI Sensor 2 Grid Bias 39188
PEMBIAS03E R-HI Sensor 3 Grid Bias 39188
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
WIFWPOSN WINDI Filter Wheel Position 39110
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 13 November 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
39107 311/06:18 (07 Nov 98) from 00.0 to 20.6 msec/day
39128 312/14:57 (08 Nov 98) from 20.6 to 00.0 msec/day
39171 315/12:48 (11 Nov 98) from 00.0 to 20.6 msec/day
39204 317/18:02 (13 Nov 98) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 13 November 1998
06 Nov 1998 - GMT Day 310 Beta = 15.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +100.8/+78.4 -3.05/-4.94 +0.0/ +0.0 16.8 0.000 0.0
2 +39.2/-11.2 +5.79/+3.87 +31.6/-14.4 24.5 1.021 21.9
3 +128.4/0. +4.51/+2.60 +30.8/-16.0 24.5 1.019 21.6
13 Nov 1998 - GMT Day 317 Beta = 45.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +100.8/+84.0 -3.05/-4.62 +0.0/ +0.0 16.8 0.000 0.0
2 +33.6/-16.8 +5.15/+3.87 +28.8/-15.2 25.8 1.026 20.8
3 +145.6/+11.2 +4.19/+2.92 +28.0/-15.2 25.8 1.023 20.4
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes (V/T 5)
APPENDIX E
UARS ANOMALY REPORTS
Week ending 13 November 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-100 FOT 39128 312/1450 N/A TAC 8 Crashed
98-101 FOT 39146 313/2108 20347 IP router reboot
98-102 FOT 39176 315/2022 N/A Table 33 errors
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-100 FOT 39128 312/14:50 TAC 8 CRASH
98-101 FOT 39146 312/21:08 20347 IP ROUTER CRASH
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-092 FOT Nov 98 See List 14170 GENERIC LATE ACQ
SEVNTFW46